# POGEMA: A Benchmark Platform for Cooperative Multi-Agent Navigation

### v1

@@ Appendix G Examples of Used Maps
The examples of used maps are presented in Figure 9, showing a diverse list of maps. The map types used in the POGEMA Benchmark include: Maze, with prolonged 1-cell width corridors requiring high-level cooperation; Random, easily generated maps to avoid overfitting with controllable obstacle density; MovingAI-tiles, smaller modified slices of MovingAI maps; Puzzle, small hand-crafted maps with challenging patterns necessitating agent cooperation; Warehouse, used in LifeLong MAPF research, featuring high agent density and throughput challenges; and MovingAI, benchmark maps with varying sizes and structures for single-agent pathfinding.
(a) Maze
@@
(c) MovingAI-tiles
(d) Puzzle
@@
(f) MovingAI
Figure 9: Examples of maps presented in the POGEMA Benchmark. The MovingAI map (on which the pathfinding metric was tested) is shown without grid lines and agents for clarity.
@@ Appendix H MARL Training Setup
For training MARL approaches, such as MAMBA, QMIX, QPLEX, and VDN, we used the default hyperparameters provided in the corresponding repositories [^12] https://github.com/jbr-ai-labs/mamba, and employed the PyMARL2 framework [^13] https://github.com/hijkzzz/pymarl2 to establish MARL baselines. As input, we apply preprocessing from the Follower approach, which is the current state-of-the-art for decentralized LifeLong MAPF. We attempted to add a ResNet encoder, as used in the Follower approach; however, this addition worsened the results, thus we opted for vectorized observation and default MLP architectures. For centralized methods that work with the state of the environment (e.g., QMIX or QPLEX), we utilized the MARL integration of POGEMA, which provides agent positions, targets, and obstacle positions in a format similar to the SMAC environment (providing their coordinates).
Our initial experiments on training this approach with a large number of agents, similar to the Follower model, showed very low results. We adjusted the training maps to be approximately $16\times 16$, which proved to be more effective and populated them with $8$ agents. This setup shows better results. We continued training the approaches until they reached a plateau, which for most algorithms is under 1 million steps.
@@ Appendix I Resources and Statistics
Random Mazes Warehouse MovingAI-tiles Puzzles MovingAI
@@
Random Mazes Warehouse MovingAI-tiles Puzzles MovingAI
@@
Our team is committed to maintaining an open and accountable POGEMA framework. Since 2021, we have continuously improved POGEMA, including the addition of the POGEMA Toolbox and the recent introduction of POGEMA Benchmark. We ensure transparency in our operations and encourage the broader AI community to participate. Our framework includes a fast learning environment, problem instance generator, visualization toolkit, and automated benchmarking tools, all guided by a clear evaluation protocol. We have also implemented and evaluated multiple strong baselines that simplify further comparison. We practice rigorous software testing and conduct regular code reviews. We promptly address issues reported on GitHub and welcome any feedback and contributions through GitHub.

### v2

@@ Appendix G Examples of Used Maps
The examples of used maps are presented in Figure 15, showing a diverse list of maps. The map types used in the POGEMA Benchmark include: Mazes, with prolonged 1-cell width corridors requiring high-level cooperation; Random, easily generated maps to avoid overfitting with controllable obstacle density; Cities-tiles, smaller modified slices of Cities maps; Puzzles, small hand-crafted maps with challenging patterns necessitating agent cooperation; Warehouse, widely used in LifeLong MAPF research, featuring high agent density and throughput challenges; and Cities, large maps with varying structures for single-agent pathfinding.
(a) Mazes
@@
(c) Cities-tiles
(d) Puzzles
@@
(f) Cities
Figure 15: Examples of maps presented in the POGEMA Benchmark. The city map (on which the pathfinding metric was tested) is shown without grid lines and agents for clarity.
@@ Appendix H MARL Training Setup
Hyperparameter IQL QPLEX QMIX VDN
Batch Size 64 64 32 32
Learning Rate 0.002 0.002 0.001 0.001
RNN Size 128 128 256 256
Table 9: Best hyperparameters found by hyperparameter sweep, which is different from defaults ones.
For training MARL approaches such as QMIX, QPLEX, IQL, and VDN, we started based on the default hyperparameters provided in the corresponding repositories, employing the PyMARL2 framework [^17] https://github.com/hijkzzz/pymarl2. These hyperparameters are mostly tuned for the SMAC environment, so we tuned the main ones for our use case. For this hyperparameter sweep, we used grid search over parameters such as learning rate, batch size, replay buffer size, and neural network parameters like the size of RNN blocks. We used the default functionality of the Wandb framework [^18] https://github.com/wandb/wandb for this sweep, with the optimization target being the CSR of the agent on the training maps. The best found hyperparameters which is different from default ones are presented in Table 9. We used the default hyperparameters for MAMBA, provided in corresponding repository [^19] https://github.com/jbr-ai-labs/mamba.
As input, we apply preprocessing from the Follower approach, which is the current state-of-the-art for decentralized LifeLong MAPF. We attempted to add a ResNet encoder, as used in the Follower approach; however, this addition worsened the results, thus we opted for vectorized observation and default MLP architectures. For centralized methods that work with the state of the environment (e.g., QMIX or QPLEX), we utilized the MARL integration of POGEMA, which provides agent positions, targets, and obstacle positions in a format similar to the SMAC environment (providing their coordinates).
Our initial experiments on training this approach with a large number of agents, similar to the Follower model, showed very low results. We adjusted the training maps to be approximately $16\times 16$, which proved to be more effective and populated them with $8$ agents. All the MARL approaches were trained using the Mazes map generator. This setup produced better results. We continued training the approaches until they reached a plateau, which for most algorithms is under 1 million steps.
@@ Appendix I Resources and Statistics
Random Mazes Warehouse Cities-tiles Puzzles Cities
@@
Random Mazes Warehouse Cities-tiles Puzzles Cities
@@
Our team is committed to maintaining an open and accountable POGEMA framework. We ensure transparency in our operations and encourage the broader AI community to participate. Our framework includes a fast learning environment, problem instance generator, visualization toolkit, and automated benchmarking tools, all guided by a clear evaluation protocol. We have also implemented and evaluated multiple strong baselines that simplify further comparison. We practice rigorous software testing and conduct regular code reviews.
Appendix K Ingestion of MovingAI Maps
We incorporated an ingestion script to convert MovingAI maps to be compatible with POGEMA. The ingestion script is straightforward, and an example of its usage is presented in Figure 16. This script downloads the full archive of, in this case, the street-map series, which will be saved to a YAML file compatible with POGEMA.
from pogema_toolbox.generators.generator_utils import ( maps_dict_to_yaml) from pogema_toolbox.moving_ai_ingestion import ( download_moving_ai_maps) url = 'https://movingai.com/benchmarks/street/street-map.zip' maps = download_moving_ai_maps(url) maps_dict_to_yaml('maps.yaml', maps)
Figure 16: Ingestion script to convert MovingAI maps into POGEMA-compatible format
Instead of converting and distributing these maps ourselves, we provide this script to allow users to convert the maps on their own, ensuring compliance with licensing terms. Additionally, by using this script, users can always work with the most up-to-date versions of the maps in the MovingAI dataset, addressing any changes or updates made to the dataset over time.
Appendix L POGEMA Speed Performance Evaluation
The speed of an environment is a critical aspect in reinforcement learning (RL), significantly influencing training performance and usability. Table 1 provides an overview of the speed performance of various multi-agent environments, demonstrating that POGEMA can process more than 10K steps per second.
Here we provide more detailed information about POGEMA’s speed performance. POGEMA’s continuous integration (CI) includes a speed measurement procedure that runs alongside the tests. Here, we present the results from a recent CI run. We compared POGEMA’s performance on three CPU setups: the AMD Ryzen Threadripper 3970X 32-Core Processor, a modern high-performance CPU; an older server-side Intel(R) Xeon(R) CPU @ 2.20GHz, commonly used on the Google Colab platform, representing an average-performance setup; and the Apple Silicon M1, a laptop setup that is widely regarded for its power efficiency and solid performance in computational tasks.
Ryzen Threadripper Intel(R) 2.20GHz Apple Silicon M1
Agents Size MAPF LMAPF MAPF LMAPF MAPF LMAPF
1 32 20 6842068420\,68420 684 21 8102181021\,81021 810 $9391$ 13 3581335813\,35813 358 29 6992969929\,69929 699 44 2244422444\,22444 224
1 64 10 3901039010\,39010 390 $9602$ $6981$ $6452$ 19 2441924419\,24419 244 18 9961899618\,99618 996
32 32 96 9189691896\,91896 918 90 6319063190\,63190 631 61 1326113261\,13261 132 61 2326123261\,23261 232 204 637204637204\,637204 637 191 069191069191\,069191 069
32 64 89 9848998489\,98489 984 85 7418574185\,74185 741 61 2976129761\,29761 297 38 3263832638\,32638 326 144 125144125144\,125144 125 175 962175962175\,962175 962
64 32 111 976111976111\,976111 976 105 558105558105\,558105 558 69 1216912169\,12169 121 39 3083930839\,30839 308 189 624189624189\,624189 624 217 482217482217\,482217 482
64 64 102 104102104102\,104102 104 96 7099670996\,70996 709 68 1266812668\,12668 126 49 9024990249\,90249 902 196 808196808196\,808196 808 194 213194213194\,213194 213
Table 12: POGEMA performance in observations per second (OPS) across different CPU types using a single CPU core. Note that the reported OPS represents the total frames (observations) received by all agents, not the environment steps.
For this test, we used default observation parameters commonly employed in learnable multi-agent pathfinding (MAPF) approaches (e.g., Follower, SCRIMP, DCC). Specifically, we set the observation radius to 5, corresponding to an $11\times 1$ 11 field. The results for both MAPF and LMAPF scenarios, using a random policy, are reported in Table 12. In scenarios with more than 32 agents, POGEMA achieves $\geq 80$K observations per second (OPS) on fast CPUs like the Ryzen Threadripper and Apple Silicon M1, and $\geq 38$K OPS on the Intel Xeon setup, which is notably fast for a single-environment, single-thread configuration. For comparison, EnvPool [^20] https://github.com/sail-sg/envpool reports 50K frames per second (FPS) for Atari on a 12-CPU setup.
We further investigate POGEMA’s speed performance using SampleFactory [^21] https://github.com/alex-petrenko/sample-factory as the sampler for parallel asynchronous execution of the environment. We used an almost default configuration with a random policy (instead of PPO) and employed two environments per worker (for double buffering in SampleFactory). The number of sampling workers varied alongside the number of agents in the environment, and the tests were run on a workstation equipped with a single AMD Ryzen Threadripper 3970X 32-Core Processor (64 threads). We conducted a sampling procedure for 5 minutes for each setup using Random maps with a size of $32\times 32$. The results both MAPF and Lifelong MAPF scenarios are presented in Figure 17.
Figure 17: Observations per second performance of POGEMA for (a) MAPF and (b) Lifelong MAPF across different number of workers and agents in the environment, using AMD Ryzen Threadripper 3970X 32-Core Processor (64 threads). We tested Random maps with size $32\times 32$.
Looking at the single-worker setup, it is notable that the performance from the previous experiment with a single CPU setup is noticeably higher. This slowdown can be attributed to the overhead produced by the parallel asynchronous execution of the framework. However, it allows to significantly improve performance; the results for the best configuration with 64 agents and 64 workers achieved 3.1M OPS for MAPF and 2.7M OPS for Lifelong MAPF. For both setups, only 16 workers are needed to exceed the significant threshold of 1M OPS. The best performance is observed in the MAPF scenario, where 916,464,000 samples were generated in 5 minutes. This demonstrates the efficiency of utilizing a high number of CPU cores for large-scale sampling tasks, which scaled almost linearly up to 32 workers. This also indicates that performance can be further improved with additional CPU resources.
To compare, we can refer to the JaxMARL paper, which provides insights into the speed performance of XLA-accelerated environments. The repository includes several environments, and we will focus on the SPS of the STORM environment, which offers grid-based tasks. Based on the paper, the speed of the environment is 2.48k with a single environment, 175k for 100 environments, and 14.6M SPS for 10,000 environments. These results were obtained using a single NVIDIA A100 GPU.
While, as expected, POGEMA is slower in very large vectorized setups compared to XLA vectorized environments that use GPU or TPU hardware acceleration, this trade-off has its advantages. It is challenging to devise an approach that can effectively utilize such a large amount of data. Additionally, by not relying on GPUs or TPUs for environment simulation, these resources remain fully available for training neural networks, which often represent the primary bottleneck in large-scale RL experiments.
We further evaluate the scalability of POGEMA with a large population of agents. For this purpose, we use a random map scenario of size $3072\times 3072$ and test agent counts starting from 1,000 agents. The results, presented in Table 13, demonstrate that POGEMA efficiently supports up to 1 million agents within a single environment.
Agents OPS SPS Reset (seconds)
1,000,000 68700.6 0.0687 173.8
100,000 67104.9 0.6710 139.6
10,000 45894.7 4.589 132.8
1,000 56477.7 56.477 120.6
Table 13: Speed performance of POGEMA with large agent populations using a random policy.
In previous experiments, we relied on a random policy to evaluate the environment’s speed performance, as it is a common choice for such tests. However, the speed performance of the environment may vary when a more advanced approach is used, as it explores a larger portion of the state space compared to the random policy. To better test the speed performance of POGEMA under real inference, we used the Follower approach and LifeLong Mazes scenarios, with a size of $128\times 128$ and up to 2048 agents, as they required creating new goals for agents upon reaching them. The results are presented in Figure 18. Here, one can see that with an increasing number of agents, the SPS of POGEMA decreases almost linearly. Additionally, the OPS throughput grows with the number of agents. This setup also highlights the ability of POGEMA to handle a large population of agents operating in the same environment. The experiment was conducted on a setup with an AMD Ryzen Threadripper 3970X 32-core processor, using a single CPU core and a single environment (no parallelization).
Figure 18: Steps per second and observations per second throughput of POGEMA with a large agent population using the Follower approach (without parallelization). Both axes are on a log scale.